# Thermodynamics of quantum processes: An operational framework for free energy and reversible athermality

### v1

@@ I.2 Outline
The paper is organized as follows. We introduce standard notations, definitions, and facts in Section II that are used to derive main results and discussion surrounding them. In Section III, we employ axiomatic approach to define the resource-theoretic free energy functions of a quantum channel and provide reasoning behind axioms. We discuss some notable definitions, namely, the (resource-theoretic) free energy, max-free energy, and hypothesis-testing free energy of quantum channels and prove some of their important properties. We introduce the resource theory of athermality for square quantum channels in Section IV and prove that the resource theory of athermality under Gibbs preserving superchannels is asymptotically reversible. We inspect quantum transmission capabilities of quantum channels and its free energy in Section V. We discuss relation between the resource theory of athermality with private randomness distillation and dynamical resource theory of purity. In Section VI, we delve deeper into the relations between the resource-theoretic free energy, thermodynamic free energy, entropy, and energy of quantum channels. We discuss operational meaning of the resource-theoretic free energy of a quantum channel with its thermodynamic utility for the maximal extractable work. We finally conclude in Section VII. The detailed proofs of several results are discussed in Appendix.
@@ II Preliminaries
The generalized entropy function $\mathbf{S}[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is defined as $\mathbf{S}[\mathcal{N}]:=-\mathbf{D}[\mathcal{N}\|\mathcal{R}^{\mathbbm{1}}]$ [24, 25], where a completely positive map $\mathcal{R}^{\omega}_{A^{\prime}\to A}$ denotes the replacer (or preparation) map that outputs $\omega\in\operatorname{Pos}(A)$, $\mathcal{R}^{\omega}_{A^{\prime}\to A}(\rho_{A^{\prime}})=\operatorname{tr}[\rho_{A^{\prime}}]\omega_{A}$ for all $\rho_{A^{\prime}}\in\operatorname{Pos}(A^{\prime})$. For a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ with $|A|<\infty$, $S_{\alpha}[\mathcal{N}]:=\ln|A|-D_{\alpha}[\mathcal{N}\|\mathcal{R}^{\pi}]$ [24], where $\alpha\in[\frac{1}{2},\infty]$ and $\alpha=1$ is $\lim_{\alpha\to 1}$. The (von Neumann) entropy of a quantum state $\rho_{A}$ is $S(\rho_{A}):=S(A)_{\rho}:=-D(\rho\|\mathbbm{1})=-\operatorname{tr}[\rho\ln\rho]$. $\mathcal{R}^{\pi}$ is called uniformly mixing or completely depolarizing channel as it always outputs the maximally mixed state $\pi(\propto\mathbbm{1})$.
@@ III Dynamical free energy
The dynamical free energy, i.e., the free energy of quantum channels, should satisfy certain desirable properties analogous to the free energy of quantum states. It becomes operationally meaningful if it is related to the athermality distillation or formation tasks under relevant free operations. It is also preferable that the channel fee energy finds operational meaning in a thermodynamic task like work extraction.
@@
    Monotonically nonincreasing under the action of an Gibbs preserving superchannel $\Theta^{\beta}$, $\Theta^{\beta}(\mathcal{T}^{\beta}_{A^{\prime}\to A})=\mathcal{T}^{\beta}_{B^{\prime}\to B}$, $f[\Theta^{\beta}(\mathcal{N})]\leq f[\mathcal{N}]$.
@@
    Convexity, $f[p\mathcal{N}+(1-p)\mathcal{M}]\leq pf[\mathcal{N}]+(1-p)f[\mathcal{M}]$ for $p\in[0,1]$, where $\mathcal{N}_{A^{\prime}\to A}$ and $\mathcal{M}_{A^{\prime}\to A}$ are quantum channels.
@@
To construct a resource-theoretic framework and coin the dynamical free energy, it is important that the free channel is the absolutely thermal channel with its output in thermal equilirbium with bath. That is, we need to fix the inverse temperature corresponding to the bath. The free energy function of a quantum channel should capture how distinguishable it is from the absolutely thermal channel [31], and should therefore be monotonically nonincreasing under the actions of Gibbs preserving superchannels. Moreover, the thermodynamic resource theory of quantum states should be recoverable from the thermodynamic resource theory of quantum channels, since quantum states can be seen as preparation channels. For the free energy function of quantum channels to be physically meaningful, it should satisfy essential properties such as continuity and additivity, and preferably convexity (or quasi-convexity), which are well-known for free energy functions of states. The axiomatic properties (A1–A6) are designed to capture these requirements.
@@
For a fixed $\beta$, the (resource-theoretic) generalized free energy function $\mathbf{F}^{\beta}[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is defined as the generalized divergence between the channel $\mathcal{N}$ and the absolutely thermal channel $\mathcal{T}^{\beta}$,
@@
For a quantum channel $\mathcal{N}_{A^{\prime}\to A}$, (a) the free energy $F^{\beta}[\mathcal{N}]:=\beta^{-1}D[\mathcal{N}\|\mathcal{T}^{\beta}]$, (b) the max-free energy $F^{\beta}_{\infty}[\mathcal{N}]:=\beta^{-1}D_{\infty}[\mathcal{N}\|\mathcal{T}^{\beta}]$, (c) the sandwiched Rényi free energy $F^{\beta}_{\alpha}[\mathcal{N}]=\beta^{-1}D_{\alpha}[\mathcal{N}\|\mathcal{T}^{\beta}]$, (d) the $\varepsilon$-hypothesis-testing free energy ${F}^{\beta,\varepsilon}_{\mathrm{H}}[\mathcal{N}]:=\beta^{-1}D^{\varepsilon}_{\mathrm{H}}[\mathcal{N}\|\mathcal{T}^{\beta}]$ for $\varepsilon\in[0,1]$, (e) ${F}^{\beta,\varepsilon}_{\infty}[\mathcal{N}]:=\beta^{-1}\inf_{\mathcal{M}\in\mathcal{B}^{\varepsilon}[\mathcal{N}]}D^{\varepsilon}_{\infty}[\mathcal{M}\|\mathcal{T}^{\beta}]$ for $\varepsilon\in[0,1]$.
@@
Proposition 2.
For total Hamiltonian of the reference to the channel and its output $\widehat{H}_{RA}=\mathbbm{1}_{R}\otimes(\widehat{H}_{A}+c\mathbbm{1}_{A})$, where $c\in\mathbbm{R}$, the max-free energy of the channel is equal to the max-free energy of its Choi state,
Proof.
The proof of Eq. (9) follows from the fact that $D_{\infty}[\mathcal{N}\|\mathcal{M}]=D_{\infty}(\Phi^{\mathcal{N}}_{RA}\|\Phi^{\mathcal{M}}_{RA})$ [48],
The proof of Eq. (10) as follows from the above. We notice that for $\widehat{H}_{RA}=\mathbbm{1}_{R}\otimes(\widehat{H}_{A}+c\mathbbm{1}_{A})$ (or, $\widehat{H}_{R}=c\mathbbm{1}_{R}$), $\pi_{R}\otimes\gamma^{\beta}_{A}=\gamma^{\beta}_{R}\otimes\gamma^{\beta}_{A}=\gamma^{\beta}_{RA}$. Then,
@@ III.1 Properties of the dynamical free energy
All the generalized free energy functions $\mathbf{F}^{\beta}[\mathcal{N}]$ satisfy (A1) and (A2) readily, whenever $\beta^{-1}\mathbf{D}(\rho\|\gamma^{\beta})$ represents some resource-theoretic free energy function $f(\rho)$ for states $\rho$ [12, 23]. The generalized free energy $\mathbf{F}^{\beta}[\mathcal{N}]$ remains invariant under the action of Gibbs preserving unitary superchannels. Faithfulness (A3) is satisfied by the generalized free energy whose generalized channel divergence is faithful. For all $\alpha\in[\frac{1}{2},\infty)$, using $\lim_{\alpha\to 1}$ as $\alpha=1$, the sandwiched Rényi free energy $F^{\beta}_{\alpha}[\mathcal{N}]\geq 0$ for an arbitrary quantum channel $\mathcal{N}_{A^{\prime}\to A}$, and $F^{\beta}_{\alpha}[\mathcal{N}]=0$ iff $\mathcal{N}=\mathcal{T}^{\beta}$. For $\alpha\geq 1$ and $n\in\mathbbm{N}$, $F_{\alpha}^{\beta}[\mathcal{N}^{\otimes n}]=nF^{\beta}_{\alpha}[\mathcal{N}]$ if the outputs of the parallel uses of the channel $\mathcal{N}$ are noninteracting (i.e., the interaction Hamiltonian between the channel outputs is zero). These assertions follow from [30, Theorem 1] and [57, Theorem 4.2].
@@
Proof.
We provide a detailed proof in Appendix A.1. ∎
Theorem 2.
Details of the proof are in Appendix A.3.
Proposition 3.
The proof is given in Appendix A.6.
@@ IV Dynamical Resource Theory of Athermality
Golden units of athermality.— For a square quantum channel $\mathcal{N}_{A^{\prime}\to A}$, the free energy is maximum if and only if the channel is unitary. For an arbitrary pair $(\mathcal{U}_{A^{\prime}\to A},\mathcal{V}_{A^{\prime}\to A})$ of unitary channels, there always exists a GPSC $\Theta^{\beta}\in\operatorname{SCh}((A^{\prime},A),(A^{\prime},A))$ such that $\Theta^{\beta}(\mathcal{U})=\mathcal{V}$, for an instance, $\Theta^{\beta}(\mathcal{N})=\mathcal{N}\circ(\mathcal{U}^{{\dagger}}\circ\mathcal{V})$. That is, all unitary channels $(\mathcal{U},\mathcal{T}^{\beta})$ are equally resourceful, $(\operatorname{id},\mathcal{T}^{\beta})\sim(\mathcal{U},\mathcal{T}^{\beta})$. This is consistent with the free energy of all unitary channels $(\mathcal{U},\mathcal{T}^{\beta})$ being the same. The free energy of a unitary channel is least when the output Hamiltonian is fully degenerate (see Lemma 9). In other words, among all unitary resource channels $(\mathcal{U}_{m},\mathcal{T}^{\beta})$ for a fixed $m$, the unitary resource channels $(\mathcal{U}_{m},\mathcal{R}^{\pi})$ have minimum free energy. With freedom of choice for golden units and the fact that GPSCs are free operations in our athermality framework, we mark our golden unit as the identity resource channel $(\operatorname{id}_{m},\mathcal{R}^{\pi})$, where $\mathcal{R}^{\pi}_{A^{\prime}\to A}$ is the absolutely thermal channel for fully degenerate output Hamiltonian ($\widehat{H}_{A}\propto\mathbbm{1}_{A}$).
It is evident that $(\operatorname{id}_{A^{\prime}\to A},\mathcal{R}^{\pi}_{A^{\prime}\to A})$ is more resourceful than $(\operatorname{id}_{B^{\prime}\to B},\mathcal{R}^{\pi}_{B^{\prime}\to B})$ if $|A|>|B|$ [24, 31, 25], therefore, we will have to be considerate of the dimension of $\operatorname{id}$ in the golden unit. For a golden unit $(\operatorname{id}_{m},\mathcal{R}^{\pi})$, where the subscript index $m$ is to denote that $\operatorname{id}$ acts on $m$-dimensional Hilbert space. The identity channel $\operatorname{id}_{m}$ (or any equivalent unitary channel $\mathcal{U}_{m}$) allows for the perfect transmission of $m^{2}$-dimensional quantum states [58, 59], equivalent to the distribution of a maximally entangled state $\Phi$ of Schmidt rank $m$ [60, 59, 27].
@@
This implies (cf. [61, Proposition 11])
$\frac{1}{2}\norm{\operatorname{id}_{m}-\mathcal{R}^{\pi}}_{\diamond}=1-\frac{1}{m^{2}}.$ (22)
@@
For an error $\varepsilon\in[0,1]$, the single-shot athermality distillation of the resource channel $(\mathcal{N},R^{\gamma})$ under GPSCs is defined as
@@
For an error $\varepsilon\in[0,1]$, the single-shot athermality formation (cost) of the resource channel $(\mathcal{N},\mathcal{R}^{\gamma})$ is defined as
@@
The single-shot athermality distillation $\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ provides the largest dimensional golden unit $(\operatorname{id}_{m},\mathcal{R}^{\pi})$ that can be distilled from the resource channel $(\mathcal{N},\mathcal{T}^{\beta})$ under GPSCs, up to allowed error $\varepsilon$. We could possibly also interpret $\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ as the maximum number of elementary golden units $(\operatorname{id}_{2},\mathcal{R}^{\pi})$ distillable from the resource channel $(\mathcal{N},\mathcal{T}^{\beta})$ under GPSCs. The single-shot athermality formation $\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ provides the least dimensional golden unit $(\operatorname{id}_{m},\mathcal{R}^{\pi})$ needed to form the resource channel $(\mathcal{N},\mathcal{T}^{\beta})$ under GPSCs, up to allowed error $\varepsilon$. We could possibly also interpret $\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ as the minimum number of elementary golden units $(\operatorname{id}_{2},\mathcal{R}^{\pi})$ spent to form the resource channel $(\mathcal{N},\mathcal{T}^{\beta})$ under GPSCs.
@@
$F^{\varepsilon}_{H}[\mathcal{N}]=\frac{2}{\beta}\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ and $F^{\varepsilon}_{\infty}[\mathcal{N}]=\frac{2}{\beta}\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$.
@@
The one-shot athermality distillation and formation of a resource channel $(\mathcal{N},\mathcal{T}^{\beta})$ are associated with how well the channel $\mathcal{N}$ can be discriminated from $\mathcal{T}^{\beta}$ in an asymmetric channel discrimination task [62, 63, 64, 65], where $\mathcal{N}$ is null hypothesis and $\mathcal{T}^{\beta}$ is alternate hypothesis. Broadly speaking, there two different strategies one can apply when multiple uses of channels are allowed, namely, adaptive and nonadaptive [62, 65]. Channel discrimination under nonadaptive strategy can be thought of as a special case of discrimination task with adaptive strategy. Nonadaptive strategy allows for parallel uses of a channel with free preprocessing and postprocessing operations, while adaptive strategy allows for sequential uses of channel with successive uses of channel being interleaved with a free adaptive channel [62, 63].
It intuitively holds that the asymptotic resource distillation rate is lesser than or equal to the formation rates in a resource theory where golden units (most resourceful objects) are distilled from lesser resource objects and spent as costs to form lesser resource objects (cf. [66]). It follows from the relation between the min relative entropy and max relative entropy, $D^{0}_{\rm H}(\rho\|\sigma)\leq D_{\infty}(\rho\|\sigma)$, that the zero-error athermality distillation and formation satisfy $\mathrm{Dist}^{0}(\mathcal{N},\mathcal{T}^{\beta})\leq\mathrm{Cost}^{0}(\mathcal{N},\mathcal{T}^{\beta})$. We now provide the trade-off relation between the single-shot athermality distillation and formation in the following proposition.

### v2

@@ I.2 Outline
The paper is organized as follows. We introduce standard notations, definitions, and facts in Section II that are used to derive main results and discussions surrounding them. In Section III, we employ an axiomatic approach to define the resource-theoretic free energy functions of a quantum channel and provide reasoning behind axioms. We discuss some notable definitions, namely, the (resource-theoretic) free energy, max-free energy, and hypothesis-testing free energy of quantum channels, and prove some of their important properties. We introduce the resource theory of athermality for square quantum channels in Section IV and prove that the resource theory of athermality under Gibbs preserving superchannels is asymptotically reversible. We inspect quantum transmission capabilities of quantum channels and its free energy in Section V. We discuss the relation between the resource theory of athermality with private randomness distillation and the dynamical resource theory of purity. In Section VI, we delve deeper into the relations between the resource-theoretic free energy, thermal free energy, entropy, and energy of quantum channels. We discuss the operational meaning of the resource-theoretic free energy of a quantum channel with its thermodynamic utility for the maximal extractable work. We finally conclude in Section VII. The detailed proofs of several results are discussed in the Appendix.
@@ II Preliminaries
The generalized entropy function $\mathbf{S}[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is defined as $\mathbf{S}[\mathcal{N}]:=-\mathbf{D}[\mathcal{N}\|\mathcal{R}^{\mathbbm{1}}]$ [24, 25], where a completely positive map $\mathcal{R}^{\omega}_{A^{\prime}\to A}$ denotes the replacer (or preparation) map that outputs $\omega\in\operatorname{Pos}(A)$, $\mathcal{R}^{\omega}_{A^{\prime}\to A}(\rho_{A^{\prime}})=\operatorname{tr}[\rho_{A^{\prime}}]\omega_{A}$ for all $\rho_{A^{\prime}}\in\operatorname{Pos}(A^{\prime})$. For a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ with $|A|<\infty$, $S_{\alpha}[\mathcal{N}]:=\ln|A|-D_{\alpha}[\mathcal{N}\|\mathcal{R}^{\pi}]$ [24], where $\alpha\in[\frac{1}{2},1)\cup(1,\infty)$. We have $\lim_{\alpha\to 1}D_{\alpha}[\mathcal{N}\|\mathcal{R}^{\pi}]=D[\mathcal{N}\|\mathcal{R}^{\pi}]$. The (von Neumann) entropy of a quantum state $\rho_{A}$ is $S(\rho_{A}):=S(A)_{\rho}:=-D(\rho\|\mathbbm{1})=-\operatorname{tr}[\rho\ln\rho]$. $\mathcal{R}^{\pi}$ is called uniformly mixing or completely depolarizing channel as it always outputs the maximally mixed state $\pi(\propto\mathbbm{1})$.
@@ III Dynamical free energy
The dynamical free energy, i.e., the free energy of quantum channels, should satisfy certain desirable properties analogous to the free energy of quantum states. It becomes operationally meaningful if it is related to athermality distillation or formation tasks under relevant free operations. It is also preferable that the channel free energy finds operational meaning in a thermodynamic task like work extraction.
@@
    Monotonically nonincreasing under the action of a Gibbs preserving superchannel $\Theta^{\beta}$, $\Theta^{\beta}(\mathcal{T}^{\beta}_{A^{\prime}\to A})=\mathcal{T}^{\beta}_{B^{\prime}\to B}$, $f[\Theta^{\beta}(\mathcal{N})]\leq f[\mathcal{N}]$.
@@
    Convexity: $f[p\mathcal{N}+(1-p)\mathcal{M}]\leq pf[\mathcal{N}]+(1-p)f[\mathcal{M}]$ for $p\in[0,1]$, where $\mathcal{N}_{A^{\prime}\to A}$ and $\mathcal{M}_{A^{\prime}\to A}$ are quantum channels.
@@
To construct a resource-theoretic framework and coin the dynamical free energy, it is important that the free channel is the absolutely thermal channel with its output in thermal equilibrium with bath. That is, we need to fix the inverse temperature corresponding to the bath. The free energy function of a quantum channel should capture how distinguishable it is from the absolutely thermal channel [31], and should therefore be monotonically nonincreasing under the actions of Gibbs preserving superchannels. Moreover, the thermodynamic resource theory of quantum states should be recoverable from the thermodynamic resource theory of quantum channels, since quantum states can be seen as preparation channels. For the free energy function of quantum channels to be physically meaningful, it should satisfy essential properties such as continuity and additivity, and preferably convexity (or quasi-convexity), which are well-known for free energy functions of states. The axiomatic properties (A1–A6) are designed to capture these requirements.
@@
For a fixed $\beta$, the (resource-theoretic) generalized free energy function $\mathbf{F}^{\beta}[\mathcal{N}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ is defined as the inverse temperature times the generalized divergence of athermality $\mathbf{D}[\mathcal{N}\|\mathcal{T}^{\beta}]$ of the quantum channel $\mathcal{N}$,
@@
For a quantum channel $\mathcal{N}_{A^{\prime}\to A}$, its generalized channel divergence with respect to the absolutely thermal channel provides quantification of athermality (nonthermality) in the channel $\mathcal{N}$; we are particularly interested in (a) the free energy $F^{\beta}[\mathcal{N}]:=\beta^{-1}D[\mathcal{N}\|\mathcal{T}^{\beta}]$, (b) the max-free energy $F^{\beta}_{\infty}[\mathcal{N}]:=\beta^{-1}D_{\infty}[\mathcal{N}\|\mathcal{T}^{\beta}]$, (c) the sandwiched Rényi free energy $F^{\beta}_{\alpha}[\mathcal{N}]=\beta^{-1}D_{\alpha}[\mathcal{N}\|\mathcal{T}^{\beta}]$, (d) the $\varepsilon$-hypothesis-testing free energy ${F}^{\beta,\varepsilon}_{\mathrm{H}}[\mathcal{N}]:=\beta^{-1}D^{\varepsilon}_{\mathrm{H}}[\mathcal{N}\|\mathcal{T}^{\beta}]$ for $\varepsilon\in[0,1]$, (e) ${F}^{\beta,\varepsilon}_{\infty}[\mathcal{N}]:=\beta^{-1}\inf_{\mathcal{M}\in\mathcal{B}^{\varepsilon}[\mathcal{N}]}D^{\varepsilon}_{\infty}[\mathcal{M}\|\mathcal{T}^{\beta}]$ for $\varepsilon\in[0,1]$.
@@ III.1 Properties of the dynamical free energy
All the generalized free energy functions $\mathbf{F}^{\beta}[\mathcal{N}]$ satisfy (A1) and (A2) readily, whenever $\beta^{-1}\mathbf{D}(\rho\|\gamma^{\beta})$ represents some resource-theoretic free energy function $f(\rho)$ for states $\rho$ [12, 23]. The generalized free energy $\mathbf{F}^{\beta}[\mathcal{N}]$ remains invariant under the action of Gibbs preserving unitary superchannels. Faithfulness (A3) is satisfied by the generalized free energy whose generalized channel divergence is faithful. For all $\alpha\in[\frac{1}{2},\infty)$, using $\lim_{\alpha\to 1}$ as $\alpha=1$, the sandwiched Rényi free energy $F^{\beta}_{\alpha}[\mathcal{N}]\geq 0$ for an arbitrary quantum channel $\mathcal{N}_{A^{\prime}\to A}$, and $F^{\beta}_{\alpha}[\mathcal{N}]=0$ iff $\mathcal{N}=\mathcal{T}^{\beta}$. For $\alpha\geq 1$ and $n\in\mathbbm{N}$, $F_{\alpha}^{\beta}[\mathcal{N}^{\otimes n}]=nF^{\beta}_{\alpha}[\mathcal{N}]$ if the outputs of the parallel uses of the channel $\mathcal{N}$ are noninteracting (i.e., the interaction Hamiltonian between the channel outputs is zero). These assertions follow from [30, Theorem 1] and [58, Theorem 4.2].
@@
We provide a detailed proof in Appendix A.1.
@@
We note that the expression for the generalized divergence of athermality $\mathbf{D}[\mathcal{N}\|\mathcal{T}^{\beta}]$ of a quantum channel $\mathcal{N}_{A^{\prime}\to A}$ was stated in [25, Eqs. (67) & (70)] in relation to its thermal entropy $\mathbf{S}^{\beta}[\mathcal{N}]=-\mathbf{D}[\mathcal{N}\|\mathcal{T}^{\beta}]$. Albeit in a different context, a quantity termed as the robustness of athermality $R^{\beta}_{T}[\mathcal{N}]$ of a quantum channel, a particular example of the generalized divergence of athermality of the quantum channel, is studied in [59]. The robustness of athermality is directly related to the max-relative entropy of athermality (and max-free energy), $D_{\infty}[\mathcal{N}\|\mathcal{T}^{\beta}]=\ln(1+R^{\beta}_{T})$ (cf. [52, 49, 60, 59]).
Lemma 2.
For total Hamiltonian of the reference to the channel and its output $\widehat{H}_{RA}=\mathbbm{1}_{R}\otimes(\widehat{H}_{A}+c\mathbbm{1}_{A})$, where $c\in\mathbbm{R}$, the max-free energy of the channel is equal to the max-free energy of its Choi state (cf. [59]),
Proof.
The proof of Eq. (9) follows from the fact that $D_{\infty}[\mathcal{N}\|\mathcal{M}]=D_{\infty}(\Phi^{\mathcal{N}}_{RA}\|\Phi^{\mathcal{M}}_{RA})$ [49] and [25, Eq. (70)],
The proof of Eq. (10) as follows from the above. We notice that for $\widehat{H}_{RA}=\mathbbm{1}_{R}\otimes(\widehat{H}_{A}+c\mathbbm{1}_{A})$ (or, $\widehat{H}_{R}=c\mathbbm{1}_{R}$), $\pi_{R}\otimes\gamma^{\beta}_{A}=\gamma^{\beta}_{R}\otimes\gamma^{\beta}_{A}=\gamma^{\beta}_{RA}$. Then [59],
Lemma 3 ([59]).
If the Hamiltonian of $A$ is fully degenerate, $\widehat{H}_{A}\propto\mathbbm{1}_{A}$, then $F^{\beta}_{\infty}[\mathcal{U}]=2\beta^{-1}\ln|A|$.
For the sake of completeness, the proof is given in Appendix A.6.
Proposition 2.
Details of the proof are in Appendix A.3; the proof for $\alpha=\infty$ follows from [59, Corollary A.4.2], since $F_{\infty}^{\beta}[\mathcal{N}]=\beta^{-1}\ln(1+R^{\beta}_{T}[\mathcal{N}])$.
@@ IV Dynamical Resource Theory of Athermality
Golden units of athermality.— For a square quantum channel $\mathcal{N}_{A^{\prime}\to A}$, the free energy is maximum if and only if the channel is unitary. For an arbitrary pair $\mathcal{U}_{A^{\prime}\to A},\mathcal{V}_{A^{\prime}\to A}$ of unitary channels, there always exists a GPSC $\Theta^{\beta}\in\operatorname{SCh}((A^{\prime},A),(A^{\prime},A))$ such that $\Theta^{\beta}(\mathcal{U})=\mathcal{V}$, for an instance, $\Theta^{\beta}(\mathcal{N})=\mathcal{N}\circ(\mathcal{U}^{{\dagger}}\circ\mathcal{V})$. That is, all unitary channels $(\mathcal{U},\mathcal{T}^{\beta})$ are equally resourceful, $(\operatorname{id}_{A^{\prime}\to A},\mathcal{T}^{\beta}_{A^{\prime}\to A})\sim(\mathcal{U}_{A^{\prime}\to A},\mathcal{T}^{\beta}_{A^{\prime}\to A})$. This is consistent with the free energy of all unitary channels $(\mathcal{U},\mathcal{T}^{\beta})$ being the same. The free energy of a unitary channel is least when the output Hamiltonian is fully degenerate (see Lemma 11). With freedom of choice for golden units and the fact that GPSCs are free operations in our athermality framework, we mark our golden unit as the identity resource channel $(\operatorname{id}_{A^{\prime}\to A},\mathcal{R}^{\pi})$, where $\mathcal{R}^{\pi}_{A^{\prime}\to A}$ is the absolutely thermal channel for fully degenerate output Hamiltonian ($\widehat{H}_{A}\propto\mathbbm{1}_{A}$). We denote $(\operatorname{id}_{A^{\prime}\to A},\mathcal{R}^{\pi}_{A^{\prime}\to A})$ as $(\operatorname{id}_{m},\mathcal{R}^{\pi})$ for $|A^{\prime}|=|A|=m$.
It is evident that $(\operatorname{id}_{A^{\prime}\to A},\mathcal{R}^{\pi}_{A^{\prime}\to A})$ is more resourceful than $(\operatorname{id}_{B^{\prime}\to B},\mathcal{R}^{\pi}_{B^{\prime}\to B})$ if $|A|>|B|$ [24, 31, 25], therefore, we will have to be considerate of the dimension of $\operatorname{id}$ in the golden unit. The identity channel $\operatorname{id}_{m}$ (or any equivalent unitary channel $\mathcal{U}_{m}$) allows for the perfect distribution of $m\times m$-dimensional quantum states between two parties [61, 62], equivalent to the distribution of a maximally entangled state $\Phi$ of Schmidt rank $m$ [63, 62, 27].
@@
This implies (cf. [64, Proposition 11])
$\frac{1}{2}\norm{\operatorname{id}_{m}-\mathcal{R}^{\pi}}_{\diamond}=\frac{1}{2}\norm{\mathcal{U}-\mathcal{R}^{\pi}}_{\diamond}=1-\frac{1}{m^{2}}.$ (22)
@@
For an error $\varepsilon\in[0,1]$, the single-shot athermality distillation of a resource channel $(\mathcal{N},\mathcal{T}^{\beta})$ under GPSCs is defined as
@@
For an error $\varepsilon\in[0,1]$, the single-shot athermality formation (cost) of a resource channel $(\mathcal{N},\mathcal{T}^{\beta})$ is defined as
@@
The single-shot athermality distillation $\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ provides the largest dimensional golden unit $(\operatorname{id}_{m},\mathcal{R}^{\pi})$ that can be distilled from the resource channel $(\mathcal{N},\mathcal{T}^{\beta})$ under GPSCs, up to allowed error $\varepsilon$. We could possibly also interpret $(\ln 2)^{-1}\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ as the maximum number of elementary golden units $(\operatorname{id}_{2},\mathcal{R}^{\pi})$ distillable from the resource channel $(\mathcal{N},\mathcal{T}^{\beta})$ under GPSCs. The single-shot athermality formation $\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ provides the least dimensional golden unit $(\operatorname{id}_{m},\mathcal{R}^{\pi})$ needed to form the resource channel $(\mathcal{N},\mathcal{T}^{\beta})$ under GPSCs, up to allowed error $\varepsilon$. We could possibly also interpret $(\ln 2)^{-1}\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ as the minimum number of elementary golden units $(\operatorname{id}_{2},\mathcal{R}^{\pi})$ spent to form the resource channel $(\mathcal{N},\mathcal{T}^{\beta})$ under GPSCs.
@@
$F^{\beta,\varepsilon}_{H}[\mathcal{N}]=\frac{2}{\beta}\mathrm{Dist}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$ and $F^{\beta,\varepsilon}_{\infty}[\mathcal{N}]=\frac{2}{\beta}\mathrm{Cost}^{\varepsilon}(\mathcal{N},\mathcal{T}^{\beta})$.
@@
The one-shot athermality distillation and formation of a resource channel $(\mathcal{N},\mathcal{T}^{\beta})$ are associated with how well the channel $\mathcal{N}$ can be discriminated from $\mathcal{T}^{\beta}$ in an asymmetric channel discrimination task [65, 66, 67, 68], where $\mathcal{N}$ is null hypothesis and $\mathcal{T}^{\beta}$ is alternate hypothesis. Broadly speaking, there are two different strategies one can apply when multiple uses of channels are allowed, namely, adaptive and nonadaptive [65, 68]. Channel discrimination under nonadaptive strategy can be thought of as a special case of discrimination task with adaptive strategy. Nonadaptive strategy allows for parallel uses of a channel with free preprocessing and postprocessing operations, while adaptive strategy allows for sequential uses of channel with successive uses of channel being interleaved with a free adaptive channel [65, 66].
It intuitively holds that the asymptotic resource distillation rate is lesser than or equal to the formation rates in a resource theory where golden units (most resourceful objects) are distilled from lesser resource objects and spent as costs to form lesser resource objects (cf. [69]). It follows from the relation between the min relative entropy and max relative entropy, $D^{0}_{\rm H}(\rho\|\sigma)\leq D_{\infty}(\rho\|\sigma)$, that the zero-error athermality distillation and formation satisfy $\mathrm{Dist}^{0}(\mathcal{N},\mathcal{T}^{\beta})\leq\mathrm{Cost}^{0}(\mathcal{N},\mathcal{T}^{\beta})$. We now provide the trade-off relation between the single-shot athermality distillation and formation in the following proposition, see Appendix A.7 for the proof.